OpenStudy (anonymous):

A square and a rectangle have equal areas. if the rectangle is 36 cm by 25cm, what is the perimeter of the square?

OpenStudy (anonymous):

Strategy: 1) Find the area of rectangle 2) Since both area are equal, we can use this area to find perimeter of the square.

OpenStudy (anonymous):

can you explain more

OpenStudy (anonymous):

1) Area of rectangle is 36 cm x 26 cm = .... (Say A) 2) \(A=s^2\), so \(s=\sqrt{A}\)

OpenStudy (anonymous):

so its 936

OpenStudy (anonymous):

Then what's the perimeter?

OpenStudy (anonymous):

but theres no answer thats 936 its only 30 60 120 900